OpenStudy (anonymous):

WILL GIVE MEDAL...pretty simple question...i just dont get it 1. In a flower garden, there are 3 tulips for every 7 daisies. If there are 15 tulips, how many daisies are there? 37 33 38 35

jimthompson5910 (jim_thompson5910):

(3 tulips)/(7 daisies) = (15 tulips)/(x daisies) 3/7 = 15/x 3*x = 7*15 ... cross multiply 3x = 105 do you see how to finish up?

OpenStudy (anonymous):

um.....

jimthompson5910 (jim_thompson5910):

you have to solve 3x = 105 for x

OpenStudy (anonymous):

35?

jimthompson5910 (jim_thompson5910):

correct